OpenStudy (anonymous):

the sume of three consecutive odd integers if five less than twice the smallest of these integers. Find the smallest of these integers and find the greatest of these integers.??

OpenStudy (radar):

let 2x+1 = first odd integer let 2x+3 = 2nd odd integer let 2x+5 = 3rd odd integer Sum = 2x+1+2x+3+2x=5=6x+9 Problem stated that the sum is 5 less than twice the smallest of these integers. Sum = 2(2x+1) - 5. write the equation 6x + 9 = 2(2x+1)-5 6x+ 9 = 4x +2 -5 6x-4x=-12 2x = -12 x=-6 first odd integer =-12+1 = -11 2nd odd integer =-12+3 = -9 3rd odd integer =-12 +5= -7 The smallest of these integers is -11 The sum of these integers is -27 which is 5 less than twice -11 or-22